OpenStudy (anonymous):

Assume f(x) = -2x + 8 and g(x) = 5x. What is the value of (f - g)(2)?

OpenStudy (anonymous):

first take f(x) - g(x) = -2x+8 -5x = -7x +8 now they ask for the value at x=2 =-7(2)+8 = -14+8 (f-g)(2)= -6

OpenStudy (anonymous):

so the answer is -6?

OpenStudy (anonymous):

Yes :o